A diabetes test is being applied in a population of 5000 men. Previous evaluation of the diabetes test in a different population showed a sensitivity of 70% and an specificity of 80%. The prevalence of diabetes is 0.5%. What is the Negative Predictive Value of the Test?
Answer #1

P(total negative tested)=P(have disease)*P(negative|have disease)+P(not have disease)*P(negative|not have disease)=0.005*(1-0.7)+(1-0.005)*0.8 =0.7975

Negative Predictive Value of the Test =true negative tested/total negative tested

=P(not have disease)*P(negative|not have disease)/P(total negative tested)

=(1-0.005)*0.8/0.7975

=0.9981